OpenStudy (anonymous):

A copper wire of length 2 m has resistance 10 ohm. What would be the resistance of the wire if the wire is stretched such that the length becomes 6 m?

OpenStudy (anonymous):

I got the new resistance as 90 ohm but am not sure of my calculation method.... Can someone help pls..............

OpenStudy (anonymous):

\[R \alpha L\] \[\Delta R \alpha \Delta L \] therefore, \[\frac{ \Delta R }{ R } = \frac{ \Delta L }{ L }\]

OpenStudy (anonymous):

@kausarsalley thanks for your help but I need to use the regular formula R=rho*l/A to prove this

OpenStudy (anonymous):

@TuringTest @.Sam. Can u pls help with this???

OpenStudy (anonymous):

if \[R=\rho \frac{ l }{ A }\] then it means that \[R \alpha L\] and if (assuming the change in cross-sectional area is negligible) \[\Delta R = \rho \frac{ \Delta l }{ A }\] then, \[\Delta R \alpha \Delta l\] but if you do not understand, its okay.....hope someone explains it better...(: